solve by substitution
4x+5y=-2
4y-x=11

Answer by checkley77(12844)

4x+5y=-2
4y-x=11
-x=-4y+11
x=4y-11
Now replace x with (4y-11) in the first equation & solve for y.
4(4y-11)+5y=-2
16y-44+5y=-2
21y=-2+44
21y=42
y=42/21
y=2 ans.
x=4*2-11
x=8-11
x=-3 ans.